OpenStudy (anonymous):

Which statement about the function below is true? f(x) = x^3 - 5x^2 - 25x + 125 = (x + 5)(x - 5)(x 5) The root 5 has a multiplicity of 1. The root 5 has a multiplicity of 2. The root 5 has a multiplicity of 1. The root 5 has a multiplicity of 3.

Directrix (directrix):

First, when you make a post, proof it. Notice that you left off a sign in your posted question: (x + 5)(x - 5) -->(x 5)

Directrix (directrix):

So, I factored the expression: If you solve this: (x - 5)(x - 5)(x + 5) = 0 By the Zero Product Property: x - 5 = 0 OR x - 5 = 0 OR x + 5 = 0 So, solve each of those linear equations for x. Then, you will know the multiplicity of the roots.

Directrix (directrix):

@bob8590